# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

### v2

@@ 4 Methodology
In this section, we first present the proposed Hierarchical Linear Quantization method. Then, we discuss its integration with existing efficient quantization techniques. Finally, we describe the memory organization strategy and LUT design that enable efficient inference on edge devices.
@@ 4.1 Hierarchical Linear Quantization
As discussed previously, most PTQ methods, such as GPTQ and AWQ, adopt uniform quantization to facilitate hardware acceleration. However, uniform quantization cannot effectively represent the distribution of weights. Motivated by the empirical observation that weight distributions in LLMs tend to follow a bell-shaped curve (Kim et al., 2023), inspired by Binary Coding (Xu et al., 2018), we propose Hierarchical Linear Quantization as an alternative approach. Specifically, for an $n$-dimensional weight vector $\mathrm{W}$, its $q$-bit quantized representation is denoted as $\mathrm{\hat{W}}$:
@@
It is worth noting that HLQ neither requires weight reconstruction nor introduces any additional factors, making it highly generalizable. It can be seamlessly integrated with various quantization methods. As a form of non-uniform quantization, it not only offers significant advantages in reducing weight representation error (see Section 4.2), but also introduces no additional computational overhead to Bit-serial LUT-based GEMM (see Section 9).
4.2 Weight Error Reduction via Hierarchical Linear Quantization
In this subsection, we introduce how to use Hierarchical Linear Quantization to reduce weight quantization error. Let $\mathrm{\hat{W}}=\mathrm{HLQ}(\mathrm{W};s,z)$, The objective is to find the optimal set of $q$ scales and a zero-point $z$ to represent $\mathrm{W}$. The optimization objective can be formulated as:
[…]
For this optimization problem, there are two possible approaches: a heuristic alternating optimization method and a gradient-based search method. The alternating optimization method features a simple computation flow and can efficiently obtain a locally convergent solution. In contrast, the gradient-based approach relies on backpropagation, which is computationally more expensive and is sensitive to the learning rates. Here, we just introduce Alternating optimization method. Gradient-based search method can be seen in Appendix A
Alternating optimization is a heuristic algorithm. At initialization, we set $z=\min(\mathrm{W})$, and the initial value of s is set to the scale factor used in uniform quantization. Specifically, let $\Delta=\frac{\max(\mathrm{W})-\min(\mathrm{W})}{2^{q}-1}$, For a $q$-bit quantization, the initial value of s is set to $s=[\Delta,2\Delta,...2^{q-1}\Delta]$. After initialization, we solve the minimization problem in formulation 7 through an alternating optimization procedure between Bit-Pattern Selection and Linear Reconstruction.
$(s^{(t)},z^{(t)})=\mathrm{LSE}{(\mathrm{W},B^{(t)})},$ (10)
where $\mathrm{LSE}$ denotes the standard least squares estimation.
We summarize alternating optimization in Algorithm 1.
@@ 4.2 Weight Error Reduction via Hierarchical Linear Quantization
0: Weight matrix $\mathrm{W}\in R^{n\times k}$, bit width q, Group size g, Max iterations $T_{max}$
@@
Figure 4 shows that HLQ significantly reduces the quantization error compared to uniform quantization, with the MSE of uniform quantization at the 1e-4 level, while HLQ decreases it to the 1e-6 level. In Table 4, we present an ablation study showing that simply replacing RTN with HLQ, without any calibration-based fine-tuning, can substantially improve model accuracy in both the 3-bit and 2-bit settings.
4.3 Post-Training Quantization for HLQ
Here, using GPTQ as an example, we illustrate how HLQ can be employed to enhance existing post-training quantization methods. GPTQ, a widely adopted PTQ approach, offers strong hardware efficiency due to its use of group-wise uniform quantization and can quantize models within a very short time. However, it suffers from significant accuracy degradation under 2-bit quantization. By integrating HLQ into the GPTQ quantization pipeline, we effectively improve its performance in both 2-bit and 3-bit settings.
Figure 5 illustrates the integration of HLQ into the GPTQ pipeline. As shown in Figure 5(a), the standard GPTQ process partitions the weights into multiple column blocks and recursively quantizes each block column by column, using unquantized columns to compensate for quantization errors. After a block is fully quantized, subsequent unquantized blocks are leveraged to further correct accumulated errors. In contrast, Figure 5(b) presents our HLQ-GPTQ procedure. Instead of recursive column-wise quantization, our method directly applies HLQ to each column block as a whole, while retaining the error compensation mechanism through subsequent blocks. In this PTQ pipeline, we only replace the block-wise quantization step with HLQ, keeping all other components identical to GPTQ. Despite this straightforward modification, it substantially enhances GPTQ’s performance in low-bit quantization scenarios (see Section 5.1.4).
Although integrating HLQ introduces additional parameter search overhead to GPTQ, this cost remains within a controllable range. Specifically, HLQ-GPTQ can still quantize the Llama2-7B model within half an hour using only 8 GB of GPU memory, thereby retaining the time and memory efficiency characteristic of PTQ methods (see Section 5.3). It is worth noting that under 2-bit quantization, the performance of HLQ-GPTQ still lags behind state-of-the-art PTQ methods such as DB-LLM (Chen et al., 2024b). However, this performance gap primarily arises from fundamental differences in methodological complexity. Methods like DB-LLM typically incorporate additional procedures such as knowledge distillation, which require significantly more computational and memory resources and often take tens of times longer to complete quantization. In contrast, HLQ-GPTQ preserves the simplicity and hardware friendliness of the PTQ paradigm. More importantly, HLQ serves as a fundamental quantization format that is largely orthogonal to most post-training optimization techniques. This implies that by integrating existing PTQ refinements, the performance of HLQ under low-bit settings can be further improved.
Figure 5: Standard GPTQ VS HLQ-GPTQ.
@@ 4.4 Efficient Finetuning for HLQ
Previous quantization methods (Shao et al., 2023; Chen et al., 2024b; Li et al., 2021; Chen et al., 2024c) typically improve quantization performance by performing block-wise search for optimal quantization parameters or end-to-end finetuning. However, most of these methods are designed for uniform quantization and are not directly applicable to HLQ. As shown in figure  6, We propose an efficient finetuning scheme tailored for HLQ, which consists of two parts: block-wise reconstruction and end-to-end tuning.
@@
where $\mathcal{F}$ denotes the mapping function of a Transformer block, $\mathbf{W}$ and $\mathbf{X}$ are the full-precision weights and activations, and $\mathrm{HLQ}$ represents the Hierarchical Linear Quantization function. To jointly consider local and block-level errors, we divide the block-wise reconstruction into two stages. In the first stage, we search for the optimal $\mathbf{W}_{int},s,z$ for each linear layer within the block using the technique described in Algorithm 1. In the second stage, we fix $\mathbf{W}_{int}$ and only optimize the scale $s$ and zero-point $z$. For each linear layer, this two-strategy preserves some local information while leveraging block-level context to further refine the parameters. Additionally, this two-stage scheme is also computationally efficient. In the first stage, parameter initialization for each linear layer is independent, allowing for batch-wise optimization. In the second stage, with $\mathbf{W}_{int}$ fixed, there is no need to access the original full-precision weights, further reducing computation and memory footprint. The experiments in Section 5.3 demonstrate the efficiency of this scheme.
After completing the block-wise reconstruction, we introduce end-to-end tuning, a training-based refinement approach. In this stage, we perform end-to-end optimization on the model using a calibration dataset. This technique is also adopted in EfficientQAT (Chen et al., 2024c), and following their practice, we update only the quantization scale $s$ during training. To avoid overfitting, we limit the tuning process to only 1–2 epochs.
To reduce memory consumption on both the GPU and CPU, we implemented strategies including lazy loading, CPU offloading, and disk offloading (see Appendix F for details). As a result, ELUTQ can quantize the LLaMA 3.1-70B model on a system configured with 64 GB of CPU RAM and two NVIDIA RTX 4090 GPUs (48 GB VRAM).
@@
4.5 Deploy with Edge Framework
In this subsection, we present the core design of the C++ inference framework tailored for Hierarchical Linear Quantization.
4.5.1 Weight-Rearrange.
4.5.2 Mirror Storage.
Given the formulation:
$\hat{\mathrm{W}}=\sum_{j=0}^{q-1}s_{j}\cdot b_{j}+z,b_{j}\in\{0,1\}^{n},$ (12)
4.5.3 Table Quantization.
@@ 5 Experiments
Our experiments consist of three parts. First, we evaluate the performance of the proposed hierarchical linear quantization method under both PTQ and efficient finetuning settings. Second, we assess the time and memory efficiency of our quantization approach. Finally, we evaluate our CPU kernel and the end-to-end inference performance of the quantized model on edge hardware.
@@ 5.1 Main Results
5.1.1 Models, Datasets and Baseline.
We conduct experiments on several widely used models, including LLaMA2 (Touvron et al., 2023), LLaMA3.1 (Dubey et al., 2024), Qwen3 (Yang et al., 2025). We use the C4 (Raffel et al., 2020) dataset as the calibration set. For 3-bit quantization, Our results are compared with weight-only methods employing uniform quantization schemes, such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024), and OmniQuant (Shao et al., 2023), ShiftAddLLM (You et al., 2024), AnyBCQ (Park et al., 2025a), GANQ (Zhao & Yuan, 2025) and SqueezeLLM (Kim et al., 2023). For 2-bit quantization, we additionally compare our method with several efficient mixed-precision quantization approaches, including PB-LLM (Shang et al., 2023), DB-LLM (Chen et al., 2024b) and ApiQ (Liao et al., 2024). In Appendix C, we also compare our method with codebook-based non-uniform quantization approaches such as Quip# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024), as well as QAT method.
5.1.2 Evaluation.
We evaluate all models using perplexity on the Wikitext2 (Merity et al., 2016) and C4 dataset. Additionally, we assess their zero-shot capabilities using the LM Harness framework (Gao et al., 2021). The evaluation tasks include ARC Easy (Clark et al., 2018), ARC Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020).
5.1.3 Configuration.
We evaluate weight-only quantization at two precision levels: INT3 and INT2, while keeping activations in full precision. For HLQ-GPTQ, following previous PTQ settings, we select 128 calibration samples from the C4 dataset. we set the maximum number of alternating optimization step $T_{max}$=10. For efficient finetuning, we randomly select 1K calibration samples from the C4 dataset. In the block-wise stage, we set the learning rate to $1\times 10^{-4}$ and train for 2 epochs, while in the end-to-end stage, the learning rate is set to $2\times 10^{-5}$ with 1 training epochs.
Table 1: A comparison of c4 perplexity ($\downarrow$) between weight-only quantization methods, with a context length of 2048. BPW represents the average number of bits per weight. Scale and zero-point are assumed to be stored in fp16 format. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. ”–” indicates that the framework does not support this model. The results on Wikitext2 can be found in Appendix G.
Method # W # G BPW L2-7 L2-13 L3.1-8 L3.1-70 Q3-8 Q3-14 Q3-32
Baseline 16 - 16 6.97 6.47 8.89 6.92 13.30 12.02 10.81
GPTQ 2 128 2.25 33.70 20.97 181.82 22.76 35.57 26.73 24.75
AWQ 2 128 2.25 1.72e5 9.41e4 2.14e6 1.57e5 2.52e6 1.28e7 5.47e7
OmniQuant 2 128 2.25 15.02 11.05 31.76 21.79 - - -
ApiQ 2 128 2.25 12.04 9.72 - - - - -
PB-LLM* - - 2.2 20.60 15.32 57.33 34.52 - - -
ShiftAddLLM 2 128 2.37 14.60 9.85 21.97 18.16 - - -
AnyBCQ 2 128 2.37 - - 21.64 - - - -
HLQ-GPTQ 2 128 2.37 14.89 9.75 25.44 19.92 24.06 18.30 14.25
HLQ-Finetuning 2 128 2.37 11.21 9.43 15.87 11.24 19.46 15.94 12.98
DB-LLM 2 64 - 9.62 8.38 19.20 - - - -
HLQ-GPTQ 2 64 2.75 13.27 9.24 20.52 17.31 20.82 16.08 13.48
HLQ-Finetuning 2 64 2.75 9.12 8.02 14.43 10.21 17.68 15.05 12.76
GPTQ 3 128 3.25 7.89 7.00 11.66 8.18 14.39 12.93 11.78
AWQ 3 128 3.25 7.84 6.94 11.58 8.32 18.51 12.91 11.92
OmniQ 3 128 3.25 7.75 6.98 11.66 8.12 - - -
SqueezeLLM(0.45%) 3 - 3.25 7.51 6.82 - - - - -
GANQ 3 - 3.25 7.51 - 10.95 - - - -
ShiftAddLLM 3 128 3.50 7.70 6.92 11.14 7.78 - - -
AnyBCQ 3 128 3.50 - - 12.23 - - - -
HLQ-GPTQ 3 128 3.50 7.74 6.90 10.80 7.96 14.15 12.79 11.49
HLQ-Finetuning 3 128 3.50 7.45 6.77 10.47 7.42 13.54 12.36 11.24
5.1.4 Perplexity Results.
Table 1 presents the perplexity results on the C4. The results on Wikitext2 can be found in Appendix G.
For HLQ-GPTQ, our method enhances the performance of GPTQ and surpasses other PTQ approaches with uniform quantization. Under the 3-bit configuration, compared to GPTQ, our approach achieves a perplexity reduction of 0.15 for the Llama2 models and 0.8 for Llama3.1-8B. In the 2-bit setting, our method substantially improves upon GPTQ, particularly for the Llama3.1-8B model, where perplexity on C4 is reduced from 181.82 to 25.44. Although 2-bit results do not surpass those of DB-LLM, we consider this reasonable. DB-LLM incorporates a complex knowledge distillation process, which demands substantial computational resources and time for finetuning. Moreover, its quantization format is less amenable to practical hardware deployment. In contrast, HLQ-GPTQ preserves the simplicity of PTQ, requiring minimal computational resources and time to complete model quantization.
@@
For HLQ-Finetuning, our method further improves model performance under low-bit quantization, particularly in the 2-bit setting. For the W2g128 configuration of Llama3.1-8B, the perplexity on C4 is further reduced to 15.87. Moreover, under the W2g64 configuration, our method surpasses DB-LLM, while requiring only 1K calibration samples compared to 20K used by DB-LLM.
5.1.5 Zero-Shot Tasks.
Table 2 presents the experimental results on five zero-shot tasks. Under 3-bit quantization, HLQ-GPTQ and HLQ-finetuning all improve the average accuracy by approximately 1% compared to GPTQ on Llama3.1-8B. For 2-bit quantization, HLQ-GPTQ achieves an improvement of roughly 9% over GPTQ on Llama3.1-8B, also outperforming AWQ and OmniQuant, while HLQ-Finetuning achieves an improvement of 20%.
Table 2: Average accuracy (%) on zero-shot tasks by lm_eval v0.4.9. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. Bold: best result; underlined: second-best. ”–” indicates that the framework does not support this model. Full results can be found in Appendix G. Acc is reported , not acc norm.
Method #W #G L2-7 L2-13 L3.1-8 L3.1-70 Q3-8 Q3-14 Q3-32B
FP16 16 - 64.13 67.81 69.25 75.33 68.10 71.29 72.15
GPTQ 3 128 61.47 66.18 64.21 71.28 63.04 67.28 66.44
AWQ 3 128 62.23 66.14 64.98 72.13 59.65 62.31 67.10
OmniQ 3 128 61.66 66.18 - - - - -
HLQ-GPTQ 3 128 62.32 66.44 65.02 72.48 65.64 69.05 67.45
HLQ-Finetuning 3 128 62.52 66.62 65.22 73.13 66.41 70.03 68.21
GPTQ 2 128 40.66 48.29 35.98 45.57 39.45 48.96 50.34
OminQ 2 128 46.98 53.56 - - - - -
PB-LLM* 2 128 37.60 38.39 - - - - -
HLQ-GPTQ 2 128 49.04 58.28 44.35 56.47 53.04 60.03 65.81
HLQ-Finetuning 2 128 55.07 59.25 57.79 66.60 59.44 64.99 67.57
Table 3: Comparison of Model Compression Rates between Uniform Quantization and Hierarchical Linear Quantization on Llama3.1-8B. Embedding and LM-head layers are excluded from quantization for all models. The symbol ’-’ denotes per-channel quantization. ”Rates($\uparrow$)” denote the compression ratio, and ”Mem($\downarrow$)” indicates the size of the compressed model. A comprehensive presentation of the results is provided in Appendix B.
Model Wbit G Uniform HLQ
@@
5.1.6 Compression Rates.
It can be observed that under the same weight precision setting, HLQ incurs a higher average bit-width compared to uniform quantization. For example, at 2-bit, the average bit-width is 2.25 for uniform quantization versus 2.37 for HLQ; at 3-bit, it is 3.25 for uniform quantization versus 3.5 for HLQ. This is because HLQ requires storing a separate scale for each bit plane. Table 3 presents the model compression rates. For LLaMA3.1-8B, HLQ at 2-bit requires an additional 0.1 GB of storage space compared to uniform quantization—an increase of approximately 5%. However, this storage overhead does not introduce additional computational cost in Bit-Serial LUT-based GEMM, enabling end-to-end inference performance on par with uniform quantization. We will provide a detailed analysis of this aspect in Section 9.
5.2 Ablation Study.
5.2.1 Ablation Study on HLQ vs. RTN Quantization
To better understand the isolated contribution of HLQ beyond standard RTN quantization, we conduct a comprehensive ablation study across multiple model scales and architectures. Specifically, we evaluate LLaMA2-7B, LLaMA2-13B, Qwen3-8B, and Qwen3-14B under both W3G128 and W2G128 settings, and only apply HLQ without any tuning technicals. As shown in Table 4, using HLQ alone yields substantial perplexity improvements under both 3-bit and 2-bit settings, demonstrating its effectiveness as a drop-in replacement for RTN.
Table 4: Comparison between HLQ and RTN quantization across multiple model under w3g128 and w2g128 configurations. Results are reported in terms of C4 perplexity, highlighting the standalone contribution of HLQ over standard RTN.
Models Configuration RTN HLQ
LLaMA2-7B w3g128 28.24 9.04
w2g128 1.10e5 450.06
LLaMA2-13B w3g128 13.37 7.52
w2g128 5.75e4 765.76
Qwen3-8B w3g128 541.95 17.60
w2g128 6.52e6 829.91
Qwen3-14B w3g128 579.17 15.43
w2g128 2.72e6 3432.17
5.2.2 Hyparameter Study for HLQ-GPTQ.

### v3

@@ 4.1 Hierarchical Linear Quantization
As discussed previously, uniform quantization cannot effectively represent the distribution of weights. Inspired by Binary Coding Xu et al. (2018), we propose Hierarchical Linear Quantization as an alternative approach. Specifically, for an $n$-dimensional weight vector $\mathrm{W}$, its $q$-bit quantized representation is denoted as $\mathrm{\hat{W}}$:
@@
Input: Weight matrix $\mathrm{W}\in R^{n\times k}$, bit width q,
Group size g, Max iterations $T_{max}$
Output: scales $s\in R^{n\times\frac{k}{g}\times q}$
zero points $z\in R^{n\times\frac{k}{g}}$
@@
4.2 Weight Error Reduction via Hierarchical Linear Quantization
Here, we introduce how to use Hierarchical Linear Quantization to reduce weight quantization error. Let $\mathrm{\hat{W}}=\mathrm{HLQ}(\mathrm{W};s,z)$, The objective is to find the optimal set of $q$ scales and a zero-point $z$ to represent $\mathrm{W}$. The optimization objective can be formulated as:
[…]
To simplify the optimization process, we propose a heuristic alternating optimization method. At initialization, we set $z=\min(\mathrm{W})$, and the initial value of s is set to the scale factor used in uniform quantization. Specifically, let $\Delta=\frac{\max(\mathrm{W})-\min(\mathrm{W})}{2^{q}-1}$, For a $q$-bit quantization, the initial value of s is set to $s=[\Delta,2\Delta,...2^{q-1}\Delta]$. After initialization, we solve the minimization problem in formulation 5 through an alternating optimization procedure between Bit-Pattern Selection and Linear Reconstruction.
$(s^{(t)},z^{(t)})=\mathrm{LSE}{(\mathrm{W},B^{(t)})},$ (8)
where $\mathrm{LSE}$ denotes the standard least squares estimation. We summarize alternating optimization in Algorithm 1.
Figure 3 shows that HLQ significantly reduces the quantization error compared to uniform quantization, with the MSE of uniform quantization at the 1e-4 level, while HLQ decreases it to the 1e-6 level. In figure 5, we present an ablation study showing that simply replacing RTN with HLQ, without any calibration-based fine-tuning, can substantially improve model accuracy in both the 3-bit and 2-bit settings.
It is worth noting that HLQ neither requires weight reconstruction nor introduces any additional factors, making it highly generalizable. It can be seamlessly integrated with various quantization methods.
@@ 4.3 Efficient Finetuning for HLQ
Previous quantization methods typically improve quantization performance by performing block-wise search for optimal quantization parameters or end-to-end finetuning. However, most of these methods are designed for uniform quantization and are not directly applicable to HLQ. As shown in figure 4, We propose an efficient finetuning scheme tailored for HLQ, which consists of two stages: block-wise reconstruction and end-to-end tuning.
@@
where $\mathcal{F}$ denotes the mapping function of a Transformer block, $\mathbf{W}$ and $\mathbf{X}$ are the full-precision weights and activations, and $\mathrm{HLQ}$ represents the Hierarchical Linear Quantization function. To jointly consider local and block-level errors, we divide the block-wise reconstruction into two sub-stages. In the first sub-stage, we search for the optimal $\mathbf{W}_{int},s,z$ for each linear layer within the block using the technique described in Algorithm 1. In the second sub-stage, we fix $\mathbf{W}_{int}$ and only optimize the scale $s$ and zero-point $z$ by minimizing the block-level output error. For each linear layer, this strategy preserves some local information while leveraging block-level context to further refine the parameters. Additionally, this strategy is also computationally efficient. In the first sub-stage, parameter initialization for each linear layer is independent, allowing for batch-wise optimization. In the second sub-stage, with $\mathbf{W}_{int}$ fixed, there is no need to access the original full-precision weights, further reducing memory footprint.
After completing the block-wise reconstruction, we introduce end-to-end tuning, a training-based refinement approach. In this stage, we perform end-to-end optimization on the model using a calibration dataset. This technique is also adopted in EfficientQAT Chen et al. (2024b), and following their practice, we update only the quantization scale $s$ during training.
To reduce memory consumption on both the GPU and CPU, we implemented strategies including lazy loading, CPU offloading, and disk offloading (see Appendix C for details). As a result, ELUTQ can quantize the LLaMA 3.1-70B model on a system configured with 64 GB of CPU RAM and 48 GB VRAM.
@@
4.4 Deployed on Edge Devices
We develop high-performance CPU and GPU kernels to support GEMM, optimize the design of the lookup tables, and enable efficient end-to-end inference. Please refer to Appendix D for more details.
@@ 5 Experiments
We evaluate our method on several advanced pre-trained models, including LLaMA2 Touvron et al. (2023), LLaMA3.1 Dubey et al. (2024), Qwen3 Yang et al. (2025). We use the C4 Raffel et al. (2020) dataset as the calibration set. Our results are compared with well-established weight-only methods, including GPTQ Frantar et al. (2022), AWQ Lin et al. (2024), and OmniQuant Shao et al. (2023), ShiftAddLLM You et al. (2024), AnyBCQ Park et al. (2025a), GANQ Zhao and Yuan (2025). For 2-bit quantization, we additionally compare our method with several efficient mixed-precision quantization approaches, including PB-LLM Shang et al. (2023), DB-LLM Chen et al. (2024a) and ApiQ Liao et al. (2024). In Appendix F, we also compare our method with codebook-based non-uniform quantization approaches such as Quip# Tseng et al. (2024) and AQLM Egiazarian et al. (2024), as well as QAT method.
We evaluate all models using perplexity on the Wikitext2 Merity et al. (2016) and C4 dataset. Additionally, we assess their zero-shot capabilities using the LM Harness framework Gao et al. (2021). The evaluation tasks include ARC Easy, ARC Challenge Clark et al. (2018), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), and PIQA Bisk et al. (2020). For finetuning, we randomly select 4K calibration samples with sequence length of 2048 from the C4 dataset. we set the maximum number of alternating optimization step $T_{max}$=10. In the block-wise stage, we set the learning rate to $1\times 10^{-4}$ and train for 2 epochs, while in the end-to-end stage, the learning rate is set to $2\times 10^{-5}$ with 1 training epochs. All the experiments are conducted on 2 RTX4090s. For inference performance, we measure the end-to-end text generation throughput on both CPUs and GPUs using a batch size of 1.
Method # W # G BPW L2-7 L2-13 L3.1-8 L3.1-70
Baseline 16 - 16 6.97 6.47 8.89 6.92
GPTQ 2 128 2.25 33.70 20.97 181.82 22.76
OmniQuant 2 128 2.25 15.02 11.05 31.76 21.79
ApiQ 2 128 2.25 12.04 9.72 22.19 -
PB-LLM* - - 2.2 20.60 15.32 57.33 34.52
ShiftAddLLM 2 128 2.37 14.60 9.85 21.97 18.16
AnyBCQ 2 128 2.37 12.17 9.85 21.64 14.15
ELUTQ 2 128 2.37 11.21 9.43 15.08 10.65
DB-LLM 2 64 - 9.62 8.38 18.74 12.87
ELUTQ 2 64 2.75 9.12 8.02 14.43 10.21
GPTQ 3 128 3.25 7.89 7.00 11.66 8.18
AWQ 3 128 3.25 7.84 6.94 11.58 8.32
OmniQ 3 128 3.25 7.75 6.98 11.66 8.12
GANQ 3 - 3.25 7.51 6.80 10.95 -
ShiftAddLLM 3 128 3.50 7.70 6.92 11.14 7.78
AnyBCQ 3 128 3.50 7.55 6.84 12.23 7.62
ELUTQ 3 128 3.50 7.45 6.77 10.47 7.42
Table 1: A comparison of c4 perplexity ($\downarrow$) between weight-only quantization methods, with a context length of 2048. BPW represents the average number of bits per weight. Scale and zero-point are assumed to be stored in fp16 format. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. ”–” indicates that the framework does not support this model. The results on Wikitext2 can be found in Appendix I.
Method #W #G L3.1-8 L3.1-70 Q3-8 Q3-14 Q3-32B
FP16 16 - 69.25 75.52 68.10 71.29 72.15
GPTQ 3 128 64.21 71.28 63.04 67.28 66.44
AWQ 3 128 64.98 72.13 59.65 62.31 67.10
OmniQ 3 128 64.46 71.55 62.76 64.35 66.89
ELUTQ 3 128 65.62 73.13 66.41 70.03 68.21
GPTQ 2 128 35.98 45.57 39.45 48.96 50.34
OminQ 2 128 40.91 53.57 45.17 53.36 55.28
PB-LLM* 2.2 - 30.51 41.46 35.18 41.67 43.79
AnyBCQ 2 128 55.21 65.38 58.32 61.74 64.81
ELUTQ 2 128 57.79 66.70 59.44 64.99 67.57
Table 2: Average accuracy (%) on five zero-shot tasks by lm_eval v0.4.9. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. Full results can be found in Appendix I. Acc is reported , not acc norm.
@@ 5.1 Main Results
Perplexity Results. Table 1 presents the perplexity results on the C4. The results on Wikitext2 can be found in Appendix I. ELUTQ further improves model performance under low-bit quantization, particularly in the 2-bit setting. For the W2g128 configuration of Llama3.1-8B, the perplexity on C4 is further reduced to 15.08. Moreover, under the W2g64 configuration, our method surpasses DB-LLM.
Zero-Shot Tasks. Table 2 reports the average zero-shot accuracy across five tasks. Under 3-bit quantization, ELUTQ consistently outperforms GPTQ, AWQ, and OmniQuant across all model scales, narrowing the gap to FP16. The advantage becomes more pronounced in the 2-bit setting, where ELUTQ achieves substantial gains over prior methods, including AnyBCQ. Compared with GPTQ, it achieves a roughly 90% improvement in average accuracy on LLaMA3.1-8B.
Compression Rates. HLQ incurs a higher average bit-width compared to uniform quantization. As shown in table 3, for example, at 2-bit, the average bit-width is 2.25 for uniform quantization versus 2.37 for HLQ. The reason is that HLQ requires storing a separate scale for each bit plane. For LLaMA3.1-8B, HLQ at 2-bit requires an additional 0.1 GB of storage space compared to uniform quantization. However, in the context of Bit-serial LUT-based GEMM, the additional storage overhead does not translate into extra computational cost. A detailed analysis can be found in Appendix H.
Model #W #G Uniform HLQ
@@
LLaMA3.1-70B 2 128 6.00 21.84 5.75 22.84
3 128 4.40 29.81 4.13 31.80
Table 3: Comparison of Model Compression Rates between Uniform Quantization and Hierarchical Linear Quantization on LLaMA3.1-8B and LLaMA3.1-70B. A comprehensive presentation of the results is provided in Appendix E.
5.2 Ablation Study
5.2.1 HLQ vs. RTN Quantization
To isolate the contribution of HLQ beyond standard RTN quantization, we do not employ any calibration-based techniques or additional optimization tricks. Instead, we simply replace RTN with HLQ. As shown in figure 5, this substitution consistently leads to substantial reductions in perplexity under both 3-bit and 2-bit settings, with particularly pronounced improvements in the 2-bit setting. These results indicate that HLQ can serve as a viable alternative to uniform quantization in low-bit regimes and may be integrated with other optimization techniques, such as QAT and knowledge distillation, to potentially improve performance at very low precision.
Figure 5: Comparison of HLQ and RTN quantization across multiple models under W3G128 and W2G128 settings. C4 perplexity ($\downarrow$) is reported, isolating the performance gain introduced by HLQ over standard RTN.
5.2.2 Hyparameter Study for $T_{max}$